OpenStudy (anonymous):

Medal To reupholster a large sofa, an interior designer needs 12 1/2 yd of fabric that costs $5.43 per yard and 5 3/4 yd of fabric that costs $6.94 per yard. Find the total cost of the two fabrics.

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

12 1/2 x $5.42 = (when multiplying money take off one cent). 5 3/4 x 5.42=

OpenStudy (anonymous):

3/4 = 75 and 1/2 = 50 or 6 in this case

OpenStudy (anonymous):

$68.29 and $31.16

OpenStudy (anonymous):

how do i multiply 12 1/2 by 5.43

OpenStudy (anonymous):

12.6 since half of 12 is 6

OpenStudy (anonymous):

how did you get 12.6

OpenStudy (anonymous):

ok i understand is 12.6

OpenStudy (anonymous):

12.5*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if ypu dont mind me asking what grade are you in

OpenStudy (anonymous):

12

OpenStudy (anonymous):

oh what school

OpenStudy (anonymous):

bramela

OpenStudy (anonymous):

oh if you were wondering i am in 6th grade and go to k12

OpenStudy (anonymous):

okay nice then how do i understand this stuff

OpenStudy (anonymous):

do you mean how do you not understand it

OpenStudy (anonymous):

no i mean how do i understand grade 12 stuff when u in grade 6

OpenStudy (anonymous):

I dont know I am in higher classes like even though I am in 6th grade I am doing like 7th and 8th grade work isnt bramela in Canada maybe it is diffrent there see I live in Broken Arrow,OK

OpenStudy (anonymous):

ohh ok nice

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

isnt the answer 12.5 * 5.42 = 68.75

OpenStudy (anonymous):

sorry 12.5 * 5.43= 68.78

OpenStudy (anonymous):

68.87*